# Random Projection Layers for Multidimensional Time Series Forecasting

### v2

@@ Abstract
All-Multi-Layer Perceptron (all-MLP) mixer models have been shown to be effective for time series forecasting problems. However, when such a model is applied to high-dimensional time series (e.g., the time series in a spatial-temporal dataset), its performance is likely to degrade due to overfitting issues. In this paper, we propose an all-MLP time series forecasting architecture, referred to as RPMixer. Our method leverages the ensemble-like behavior of deep neural networks, where each individual block within the network acts like a base learner in an ensemble model, especially when identity mapping residual connections are incorporated. By integrating random projection layers into our model, we increase the diversity among the blocks’ outputs, thereby enhancing the overall performance of RPMixer. Extensive experiments conducted on large-scale spatial-temporal forecasting benchmark datasets demonstrate that our proposed method outperforms alternative methods, including both spatial-temporal graph models and general forecasting models.
1 Introduction
Multidimensional time series forecasting systems play a crucial role in addressing various real-world problems Yan et al. (2018); Yao et al. (2018); Geng et al. (2019); Liu et al. (2020); Cai et al. (2023). Over the years, a multitude of multidimensional time series forecasting methods have been proposed Shang et al. (2021); Zeng et al. (2023); Chen et al. (2023). In this paper, we have a particular interest in solutions that exclusively utilize Multi-Layer Perceptron (MLP) due to its simplicity, efficiency, and state-of-the-art performance in time series forecasting Chen et al. (2023). These models, often dubbed all-MLP or mixer models, consist of layers of mixer blocks Tolstikhin et al. (2021); Chen et al. (2023). However, when applying existing methods like TSMixer Chen et al. (2023) to high-dimensional datasets, such as the multidimensional time series from spatial-temporal datasets, the performance tends to suffer, likely due to overfitting.
@@ 1 Introduction
Our proposed method seeks to enhance the existing mixer model’s capacity to capture the relationship between different dimensions of the input time series. Specifically, we incorporate random projection into mixer models to reinforce the model’s ability to learn such relationships. It has been observed that deep learning models exhibit ensemble-like behavior when identity mapping connections are integrated into the model design Veit et al. (2016). This results in residual units, or mixer blocks in the context of mixer models, functioning similarly to base learners in an ensemble model Veit et al. (2016). Given that diversity is a critical factor for the success of an ensemble model Margineantu and Dietterich (1997), and diversifying the intermediate outputs is akin to diversifying the base learners in an ensemble model, our proposed method has demonstrated superior performance compared to existing solutions on large-scale spatial-temporal forecasting benchmark datasets.
Figure 1: Each row consists of the intermediate and final outputs from a sampled node, while each column consists of the outputs from different mixer blocks within the model. The three samples are from node 238, 280, and 295 of the SD dataset (see Section 5 for details).
@@
To illustrate that the inclusion of random projection does diversify the intermediate representation, we refer to Fig. 1, which presents three examples of the outputs from the first, third, fifth, seventh, and final (eighth) blocks of two mixer models. One of these models integrates random projection (i.e., “proposed" in Fig. 1), while the other does not. As shown, the outputs of the mixer blocks in our proposed method exhibit greater diversity (i.e., the outputs from different blocks are less similar) compared to those of alternative models that lack random projection. This diversification leads to more accurate predictions, as evidenced by the output of the final block, which closely mirrors the ground truth. Due to our application of random projection in the mixer model design, we have aptly named our proposed method RPMixer.
@@
  • We have developed a novel spatial-temporal forecasting method, RPMixer, which is both effective and efficient. We incorporate random projection layers into RPMixer to enhance the modeling diversity and capability of mixer models.
  • Our ablation study highlights the importance of identity mapping for achieving the best performance of our proposed RPMixer.
  • Extensive experiments validate the effectiveness of our proposed method on large-scale spatial-temporal forecasting datasets, demonstrating superior performance compared to existing solutions.
2 Related Work
@@
[…]
@@
The spatial-temporal forecasting problem, in the absence of graph input, essentially transforms into a multidimensional time series forecasting problem (see Section 3 for details). As such, we also review methods designed for multidimensional time series forecasting. Over the years, transformer-based methods Vaswani et al. (2017) like LogTrans Li et al. (2019), Pyraformer Liu et al. (2021), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and Fedformer Zhou et al. (2022) have emerged. However, simple linear models outperform these transformer-based methods on multidimensional long-term time series forecasting datasets as Zeng et al. (2023) demonstrates. In contrast, mixer-based methods Tolstikhin et al. (2021), like TSMixer Chen et al. (2023), have demonstrated promising performance on multidimensional time series forecasting problems. Consequently, we aim to develop a mixer-based architecture for the spatial-temporal forecasting problem.
@@
Random projection Bingham and Mannila (2001) is utilized in a variety of machine learning and data mining methods Bingham and Mannila (2001); Schclar and Rokach (2009); Cannings and Samworth (2017); Yeh et al. (2022a). However, the majority of these works Bingham and Mannila (2001); Cannings and Samworth (2017); Yeh et al. (2022a) concentrate on the efficiency gains from random projection. Only Schclar and Rokach (2009) discusses how the diversity introduced by random projection could enhance the performance of an ensemble model, but without confirming that random projection actually introduces diversity into the model. Besides random projection, randomized methods like random shapelets/convolutions Renard et al. (2016); Yeh and Keogh (2016); Yeh (2018); Dempster et al. (2020) also perform exceptionally well for time series data on classification problems. Some papers also employ neural networks with partially fixed random initialized weights Huang et al. (2015); Ulyanov et al. (2018); Rosenfeld and Tsotsos (2019), but these papers neither focus on the spatial-temporal forecasting problem nor provide an analysis based on the ensemble interpretation of deep neural networks Veit et al. (2016). To the best of our knowledge, our paper is the first to investigate fixed randomized layers (or random projection) in the context of spatial-temporal forecasting problems.
3 Definition and Problem
@@
The problem formulation aligns with the multidimensional time series forecasting problem [^1] Each node in the spatial-temporal data is considered a dimension of a multidimensional time series. Zeng et al. (2023) if the adjacency matrix $A$ is disregarded by the model $F(\cdot)$. It is important to note that each node in a spatial-temporal dataset can be associated with multiple feature dimensions. This means that the input time series to the model $F(\cdot)$ could become a time series tensor $\mathbf{X}_{\text{past}}\in\mathbb{R}^{n\times d\times t_{\text{past}}}$, where $d$ represents the number of features of the time series and $t_{\text{past}}$ is the number of time steps in the input time series. Linear or MLP-based models can easily accommodate this additional dimensionality by reshaping the time series tensor $\mathbf{X}_{\text{past}}$ into a matrix with the shape $n\times dt_{\text{past}}$ as they utilize fully connected (linear) layers. Given the $dt_{\text{past}}$-sized vector associated with an entity, a fully connected layer is capable of utilizing information from every dimension in each past time step for the prediction.
4 Proposed Method
The proposed RPMixer adopts an all-MLP mixer architecture akin to the TSMixer Chen et al. (2023) method, but it differs in three aspects:
@@ 4 Proposed Method
    Our design emphasizes the inclusion of identity mapping connections, achieved with a pre-activation design. The RPMixer is inspired by the ensemble interpretation of the residual connection Veit et al. (2016), which necessitates the incorporation of identity mapping connections in the design.
@@
Figure 2: The RPMixer architecture.
@@
Figure 3: The detailed design of the mixer block. The red line highlighted the identity mapping connection for a mixer block.
The mixer block comprises two sub-blocks: 1) a temporal mixer block focusing on modeling the relationship between different time steps, and 2) a spatial mixer block focusing on modeling the relationship between different nodes. The temporal mixer block employs a complex linear layer after the ReLU activation function to model the time series in the frequency domain. We delve into the details of the complex linear layer in Appendix A.1.
@@
4.1 Random Projection Layer
@@
As demonstrated in the above pseudocode, the random projection layer is a fixed, randomly initialized linear layer that performs random projection Bingham and Mannila (2001) on the input. If we split the input time series matrix $X\in\mathbb{R}^{n\times t_{\text{past}}}$ into $t_{\text{past}}$ vectors of size $n$, the random projection layer computes $n_{\text{rand}}$ random combinations of the vector elements to form new vectors of size $n_{\text{rand}}$. In essence, by keeping the weights of the random projection layers fixed during training, we encourage each mixer block to concentrate on different nodes, since each random projection layer possesses its own set of randomly initialized weights.
Figure 4: The time series output from a random projection layer retains the daily and weekly patterns of the original time series.
According to the Johnson-Lindenstrauss lemma Johnson (1984), the relationships between the $t_{\text{past}}$ vectors are preserved after the random projection layer with high probability. This implies that the dynamics of the overall spatial-temporal data are largely retained in the data, even if it is projected to a much smaller space. To visually verify this claim, we randomly sampled nine nodes from the LargeST dataset Liu et al. (2023) to form a nine-dimensional time series and used random projection to reduce the dimensionality to three. The visualization is shown in Fig. 4, where the daily and weekly patterns are preserved in the time series output by random projection. A similar phenomenon has been observed in Yeh et al. (2023b) for the time series discord mining problem.
Another advantage of the random projection layer is its efficiency. As a dimension reduction technique Bingham and Mannila (2001), random projection can improve the model’s efficiency, especially when dealing with spatial-temporal data with a large number of nodes, by reducing the vector size from $n$ to a smaller $n_{\text{rand}}$. Overall, the random projection layers promote the diversity among each mixer block’s outputs while ensure efficiency.
4.2 Identity Mapping with Pre-Activation
The identity mapping connection He et al. (2016) plays an important role in the ensemble interpretation Veit et al. (2016) of the proposed RPMixer. It facilitates the creation of shorter paths, which acts like base learners within the model. We have implemented the identity mapping connection drawing inspiration from the pre-activation design proposed in He et al. (2016), where the activation functions (e.g., ReLU) precede the weighted layers.
@@ 4.2 Identity Mapping with Pre-Activation
$F_{\text{temp}}(X)\vcentcolon=\textsc{ComplexLinear}(\textsc{ReLU}(X))$ (1)
$F_{\text{sp}}(X)\vcentcolon=\textsc{Linear}(\textsc{ReLU}(\textsc{RandProject}(\textsc{ReLU}(X^{T}))))^{T}$ (2)
@@
$\textsc{Mixer}(X)=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)+X$ (3)
@@
$G(X)\vcentcolon=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)$ (4)
@@
$\textsc{Mixer}(X)=G(X)+X$ (5)
Figure 5: An example used for illustrating how the RPMixer could be interpreted as an ensemble model. In this example, there are only three mixer blocks in the model; however, the same analysis can be extended to models with more blocks.
@@
Figure 6: The unraveled view of different forward passes for the model illustrated in Fig. 5.
Following the analysis presented in Veit et al. (2016), we can unravel the forward pass in Fig. 5 into multiple paths. The unraveled view is illustrated in Fig. 6.
@@
5 Experiment
In this section, we present experiment results that demonstrate the effectiveness of our proposed method. We begin by introducing the dataset, benchmark settings, and baseline methods. Following this, we explore the benchmark results and an ablation study, which showcases the impact of each design choice. Importantly, we show how the random projection layer significantly enhances the diversity of the network’s intermediate representation, contributing substantially to the superior performance of our proposed method. Next, we conduct a sensitivity analysis on the crucial hyper-parameters. It is worth noting that our proposed method is a general approach for multivariate time series forecasting. Thus, we also perform experiments on benchmark datasets for multivariate time series forecasting and include the results in Appendix A.6. Details regarding the implementation can be found in Appendix A.4. The source code for our experiments can be downloaded from The Author(2023) (s).
5.1 Dataset and Benchmark Setting
We conducted our experiments using the LargeST dataset Liu et al. (2023), which consists of traffic data collected from 8,600 sensors in California from 2017 to 2021. For benchmarking purposes, we generated four sub-datasets following the procedure outlined in Liu et al. (2023). These sub-datasets, namely SD, GBA, GLA, and CA, include sensor data from the San Diego region, the Greater Bay Area region, the Greater Los Angeles region, and all 8,600 sensors across California, respectively. The statistics pertaining to these datasets are detailed in Appendix A.2.
In line with the experimental setup described in Liu et al. (2023), we only utilized traffic data from 2019. Sensor readings, originally recorded at 5-minute intervals, were aggregated into 15-minute windows, yielding 96 windows per day. Each sub-dataset was chronologically divided into training, validation, and test sets at a ratio of 6:2:2. The benchmark task was to predict the next 12 steps for each sensor at each timestamp. Performance was measured using the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE).
5.2 Baseline Method
[…]
Table 1: Performance comparisons. We bold the best-performing results. The performance reported in the “Average" column is computed by averaging over 12 predicted time steps. The absence of baselines on the GLA and CA datasets indicates that the models incur an out-of-memory issue. Param: the number of learnable parameters. K: $10^{3}$. M: $10^{6}$.

### v4

# RPMixer: Shaking Up Time Series Forecasting with Random Projections for Large Spatial-Temporal Data

@@ Abstract.
Spatial-temporal forecasting systems play a crucial role in addressing numerous real-world challenges. In this paper, we investigate the potential of addressing spatial-temporal forecasting problems using general time series forecasting models, i.e., models that do not leverage the spatial relationships among the nodes. We propose a all-Multi-Layer Perceptron (all-MLP) time series forecasting architecture called RPMixer. The all-MLP architecture was chosen due to its recent success in time series forecasting benchmarks. Furthermore, our method capitalizes on the ensemble-like behavior of deep neural networks, where each individual block within the network behaves like a base learner in an ensemble model, particularly when identity mapping residual connections are incorporated. By integrating random projection layers into our model, we increase the diversity among the blocks’ outputs, thereby improving the overall performance of the network. Extensive experiments conducted on the largest spatial-temporal forecasting benchmark datasets demonstrate that the proposed method outperforms alternative methods, including both spatial-temporal graph models and general forecasting models.
time series, forecasting, large spatial-temporal graph
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671881isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Neural networksccs: Information systems Spatial-temporal systems
1. Introduction
Spatial-temporal forecasting systems are instrumental in addressing numerous real-world problems (Yan et al., 2018; Yao et al., 2018; Geng et al., 2019; Liu et al., 2020; Cai et al., 2023). Traffic flow prediction, in particular, has attracted considerable attention due to its potential to significantly impact urban planning, traffic management, and public safety (Lan et al., 2022; Fan et al., 2023; Liu et al., 2023b). In light of this, our study specifically uses the largest traffic flow prediction benchmark dataset, LargeST (Liu et al., 2023b), to study the spatial-temporal forecasting problem.
Conventionally, graph-based methods are utilized to solve spatial-temporal forecasting problems in the literature (Liu et al., 2023b). These methods predict the future by leveraging both temporal (i.e., time series) and spatial (i.e., graph) information. However, it is important to note that these two types of information may not hold equal significance. For instance, methods like DSTAGNN (Lan et al., 2022) achieve state-of-the-art performance by relying solely on the input time series, bypassing the input graph and inferring the spatial relationship from the time series. As spatial relationships are learned, this type of method has the potential to uncover relationships that do not exist in the input graph. While these methods (Shang et al., 2021; Lan et al., 2022; Shao et al., 2022a) have shown promising performance, they often confront computational challenges when dealing with large-scale spatial-temporal datasets. This is largely due to the computation of the pairwise similarity matrix (or adjacency matrix), which has quadratic space complexity relative to the number of nodes in the dataset, and forms part of the model’s intermediate representation. This observation prompts an intriguing question: How can we maintain scalability while achieving high performance when the input graph is not utilized?
In the absence of an input graph, the spatial-temporal forecasting can be framed as a multidimensional time series forecasting problem (Zeng et al., 2023). In this context, each node is perceived as a dimension within the multidimensional time series. Over the years, numerous multidimensional time series forecasting methods have been proposed (Shang et al., 2021; Zeng et al., 2023; Chen et al., 2023). We are particularly interested in solutions that exclusively utilize Multi-Layer Perceptron (MLP), due to its simplicity, efficiency, and its state-of-the-art performance in multidimensional time series forecasting (Chen et al., 2023). These models, often referred to as all-MLP or mixer models, comprise layers of mixer blocks (Tolstikhin et al., 2021; Chen et al., 2023). Unlike graph-based methods, these models avoid the explicit computation of the computationally expensive pairwise similarity matrix.
@@ 1. Introduction
Our proposed method enhances the existing mixer model’s ability to capture both the spatial and temporal aspects of the input time series. Specifically, we incorporate random projection into mixer models to bolster the model’s capability to learn node relationships. It has been observed that when identity mapping connections are integrated into the model design, deep learning models exhibit ensemble-like behavior (Veit et al., 2016). This results in residual units, or mixer blocks in context of mixer models, functioning similarly to base learners in ensemble model (Veit et al., 2016). Given that diversity is a critical factor for the success of an ensemble model (Margineantu and Dietterich, 1997), and diversifying the intermediate outputs is akin to diversifying the base learners in an ensemble, our proposed method has demonstrated superior performance when compared to existing solutions on large-scale spatial-temporal forecasting benchmark datasets.
@@
To demonstrate the inclusion of random projection does diversify the intermediate representation, we refer to Fig. 1, which presents three examples of the outputs from the first, third, fifth, seventh, and final (eighth) blocks of two mixer models. One of these models integrates random projection (i.e., “proposed” in Fig. 1), while the other does not. As illustrated, the outputs of the mixer blocks in our proposed method exhibit greater diversity (i.e., the outputs from different blocks bear less similarity) compared to those of alternative models that lack random projection. This diversification facilitates more accurate predictions, as evidenced by the output of the final block, which closely mirrors the ground truth. Owing to our application of random projection in the mixer model design, we have aptly named our proposed method RPMixer.
Figure 1. Each row consists of the intermediate and final outputs from a sampled node, while each column consists of the outputs from different mixer blocks within the model. The three samples are from node 238, 280, and 295 of the SD dataset (see Section 5 for details).
Beyond enhancing the spatial modeling capability of mixer models, we also bolster their temporal modeling capability by processing time series in the frequency domain. This is particularly beneficial for spatial-temporal time series, which often exhibit periodicity, especially those influenced by human activity (Yan et al., 2018; Yao et al., 2018; Geng et al., 2019; Liu et al., 2020; Cai et al., 2023). Such periodicity is evident in traffic flow forecasting datasets, where daily and weekly patterns are common, as illustrated in Fig. 2. To effectively model these periodic signals, we employ the fast Fourier transform along with linear layers that handle complex numbers. This approach enables us to capture patterns and trends in the data that might otherwise be overlooked by alternative designs.
Figure 2. The time series for three randomly selected nodes from the LargeST dataset spans a period of four weeks.
@@
  • We develop a novel spatial-temporal forecasting method, RPMixer, that does not rely on the input graph and is both effective and efficient. We incorporate random projection layers in RPMixer to enhance the spatial modeling capability of mixer models.
  • We enhance the temporal modeling capability of mixer models by processing time series in the frequency domain, leveraging the inherent periodicity of many spatial-temporal datasets.
  • Extensive experiments validate the effectiveness of our proposed method on large-scale spatial-temporal forecasting datasets, demonstrating superior performance over existing solutions.
2. Related Work
@@
Spatial-temporal forecasting methods primarily use one of two spatial representations: 1) a graph-based format (Liu et al., 2023b), where the spatial information is stored in a graph, and 2) a grid-based format (Zhang et al., 2023), where nodes are arranged in a grid resembling an image. Given that the LargeST dataset (Liu et al., 2023b) is graph-based, our review focuses on graph-based methods. These methods typically integrate graph modeling architectures, such as the graph convolutional network (Kipf and Welling, 2016) or the graph attention network (Veličković et al., 2017), with sequential modeling architectures like recurrent neural networks (Hochreiter and Schmidhuber, 1997), temporal convolutional networks (Oord et al., 2016), or transformers (Vaswani et al., 2017). Recent proposals include DCRNN (Li et al., 2017), STGCN (Yu et al., 2017), ASTGCN (Guo et al., 2019), GWNET (Wu et al., 2019), AGCRN (Bai et al., 2020), STGODE (Fang et al., 2021), DSTAGNN (Lan et al., 2022), D2STGNN (Shao et al., 2022d), and DGCRN (Li et al., 2023a). We incorporate all these methods into our experiments due to their relevance to our problem. We delve into the details of these spatial-temporal forecasting methods in Appendix A.2. Note, MLPST (Zhang et al., 2023), much like RPMixer, is also a mixer-type forecasting model. However, it was specifically developed for grid-based spatial-temporal forecasting problems and, therefore, is not considered in this paper.
@@
The spatial-temporal forecasting problem, in the absence of graph input, essentially transforms into a multidimensional time series forecasting problem (see Section 3 for details). As such, we also review methods designed for multidimensional time series forecasting. Over the years, transformer-based methods (Vaswani et al., 2017) like LogTrans (Li et al., 2019), Pyraformer (Liu et al., 2021), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), and Fedformer (Zhou et al., 2022) have emerged. However, simple linear models outperform these transformer-based methods on multidimensional long-term time series forecasting datasets as (Zeng et al., 2023) demonstrates. In contrast, mixer-based methods (Tolstikhin et al., 2021), like TSMixer (Chen et al., 2023), have demonstrated promising performance on multidimensional time series forecasting problems. Consequently, we aim to develop a mixer-based architecture for the spatial-temporal forecasting problem.
@@
Random projection (Bingham and Mannila, 2001) is utilized in a variety of machine learning and data mining methods (Bingham and Mannila, 2001; Schclar and Rokach, 2009; Cannings and Samworth, 2017; Yeh et al., 2022a). However, the majority of these works (Bingham and Mannila, 2001; Cannings and Samworth, 2017; Yeh et al., 2022a) concentrate on the efficiency gains from random projection. Only (Schclar and Rokach, 2009) discusses how the diversity introduced by random projection could enhance the performance of an ensemble model, but without confirming that random projection actually introduces diversity into the model. Besides random projection, randomized methods like random shapelets/convolutions (Renard et al., 2016; Yeh and Keogh, 2016; Yeh, 2018; Dempster et al., 2020) also perform exceptionally well for time series data on classification problems. Some papers also employ neural networks with partially fixed random initialized weights (Huang et al., 2015; Ulyanov et al., 2018; Rosenfeld and Tsotsos, 2019), but these papers neither focus on the spatial-temporal forecasting problem nor provide an analysis based on the ensemble interpretation of deep neural networks (Veit et al., 2016). To the best of our knowledge, our paper is the first to investigate fixed randomized layers (or random projection) in the context of spatial-temporal forecasting problems.
3. Definition and Problem
@@
The problem formulation aligns with the multidimensional time series forecasting problem [^1] Each node in the spatial-temporal data is considered a dimension of a multidimensional time series. (Zeng et al., 2023) if the adjacency matrix $A$ is disregarded by the model $F(\cdot)$. It is important to note that each node in a spatial-temporal dataset can be associated with multiple feature dimensions. This means that the input time series to the model $F(\cdot)$ could become a time series tensor $\mathbf{X}_{\text{past}}\in\mathbb{R}^{n\times d\times t_{\text{past}}}$, where $d$ represents the number of features of the time series and $t_{\text{past}}$ is the number of time steps in the input time series. Linear or MLP-based models can easily accommodate this additional dimensionality by reshaping the time series tensor $\mathbf{X}_{\text{past}}$ into a matrix with the shape $n\times dt_{\text{past}}$ as they utilize fully connected (linear) layers. Given the $dt_{\text{past}}$-sized vector associated with an entity, a fully connected layer is capable of utilizing information from every dimension in each past time step for the prediction.
4. Proposed Method
The proposed RPMixer adopts an all-MLP mixer architecture akin to the TSMixer (Chen et al., 2023) method, but it differs in three aspects:
@@ 4. Proposed Method
    Our design emphasizes the inclusion of identity mapping connections, achieved with a pre-activation design. The RPMixer is inspired by the ensemble interpretation of the residual connection (Veit et al., 2016), which necessitates the incorporation of identity mapping connections in the design.
@@
Figure 3. The RPMixer architecture.
@@
Figure 4. The detailed design of the mixer block. The red line highlighted the identity mapping connection for a mixer block.
The mixer block comprises two sub-blocks: 1) a temporal mixer block focusing on modeling the relationship between different time steps, and 2) a spatial mixer block focusing on modeling the relationship between different nodes. The temporal mixer block employs a complex linear layer after the ReLU activation function to model the time series in the frequency domain. We delve into the details of the complex linear layer in Section 4.1.
@@
4.1. Complex Linear Layer
As depicted in Fig. 2, the time series in the large spatial-temporal forecasting benchmark dataset exhibits periodicity [^2] Please note, periodicity is an important property that enables complex linear layers to enhance performance and ensures the predictability of a time series dataset.. To leverage this periodicity, we opt to process the time series in the frequency domain using a complex linear layer (Trabelsi et al., 2018). The computational graph for the complex linear layer is illustrated in Fig. 5. Initially, the input data is converted to the frequency domain using the Fast Fourier Transform (FFT) method. Subsequently, the real and imaginary parts are processed by two distinct linear layers, one containing the real part of the model weights and the other containing the imaginary part. The outputs from these two linear layers are then combined, and finally an inverse FFT layer is used to convert back to the time domain.
Figure 5. The complex linear layer.
(1) $(W_{\text{real}}x_{\text{real}}-W_{\text{imag}}x_{\text{imag}})+{i\mkern 1.0mu}(W_{\text{real}}x_{\text{imag}}+W_{\text{imag}}x_{\text{real}})$
This operation is captured in the design illustrated in Fig. 5.
4.2. Random Projection Layer
@@
As demonstrated in the above pseudocode, the random projection layer is a fixed, randomly initialized linear layer that performs random projection (Bingham and Mannila, 2001) on the input. If we split the input time series matrix $X\in\mathbb{R}^{n\times t_{\text{past}}}$ into $t_{\text{past}}$ vectors of size $n$, the random projection layer computes $n_{\text{rand}}$ random combinations of the vector elements to form new vectors of size $n_{\text{rand}}$. In essence, by keeping the weights of a random projection layer fixed during training, we encourage the associated mixer block to concentrate on a random set of nodes. Since each random projection layer has its unique set of randomly initialized weights, different mixer blocks would focus on different sets of random nodes. Consequently, the outputs of these mixer blocks are likely to deviate from one another.
According to the Johnson-Lindenstrauss lemma (Johnson, 1984), the relationships between the $t_{\text{past}}$ vectors are preserved after the random projection layer with high probability. This implies that the dynamics of the overall spatial-temporal data are largely retained in the data, even if it is projected to a much smaller space. To visually verify this claim, we randomly sampled nine nodes from the LargeST dataset (Liu et al., 2023b) to form a nine-dimensional time series and used random projection to reduce the dimensionality to three. The visualization is shown in Fig. 6, where the daily and weekly patterns are preserved in the time series output by random projection. A similar phenomenon has been observed in (Yeh et al., 2023c) for the time series discord mining problem.
Figure 6. The time series output from a random projection layer retains the daily and weekly patterns of the original time series.
4.3. Identity Mapping with Pre-Activation
The identity mapping connection (He et al., 2016) plays an important role in the ensemble interpretation (Veit et al., 2016) of the proposed RPMixer. It facilitates the creation of shorter paths, which acts like base learners within the model. We have implemented the identity mapping connection drawing inspiration from the pre-activation design proposed in (He et al., 2016), where the activation functions (e.g., ReLU) precede the weighted layers.
@@ 4.3. Identity Mapping with Pre-Activation
(2) $F_{\text{temp}}(X)\vcentcolon=\textsc{ComplexLinear}(\textsc{ReLU}(X))$
(3) $F_{\text{sp}}(X)\vcentcolon=\textsc{Linear}(\textsc{ReLU}(\textsc{RandProject}(\textsc{ReLU}(X^{T}))))^{T}$
@@
(4) $\textsc{Mixer}(X)=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)+X$
@@
(5) $G(X)\vcentcolon=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)$
@@
(6) $\textsc{Mixer}(X)=G(X)+X$
@@
Figure 7. An example used for illustrating how the RPMixer could be interpreted as an ensemble model. In this example, there are only three mixer blocks in the model; however, the same analysis can be extended to models with more blocks.
Following the analysis presented in (Veit et al., 2016), we can unravel the forward pass in Fig. 7 into multiple paths. The unraveled view is illustrated in Fig. 8.
Figure 8. The unraveled view of different forward passes for the model illustrated in Fig. 7.
@@
5. Experiment
In this section, we present experiment results that demonstrate the effectiveness of our proposed method. We begin by introducing the dataset, benchmark settings, and baseline methods. Following this, we explore the benchmark results and an ablation study, which showcases the impact of each design choice. Importantly, we show how the random projection layer significantly enhances the diversity of the network’s intermediate representation, contributing substantially to the superior performance of our proposed method. Next, we conduct a sensitivity analysis on the crucial hyper-parameters. It is worth noting that our proposed method is a general approach for multivariate time series forecasting. Thus, we also perform experiments on benchmark datasets for multivariate time series forecasting and include the results in Appendix A.7. Details regarding the implementation can be found in Appendix A.3. The source code for our experiments can be downloaded from (The Author(2023), s).
5.1. Dataset and Benchmark Setting
We conducted our experiments using the LargeST dataset (Liu et al., 2023b), which consists of traffic data collected from 8,600 sensors in California from 2017 to 2021. For benchmarking purposes, we generated four sub-datasets following the procedure outlined in (Liu et al., 2023b). These sub-datasets, namely SD, GBA, GLA, and CA, include sensor data from the San Diego region, the Greater Bay Area region, the Greater Los Angeles region, and all 8,600 sensors across California, respectively. The statistics pertaining to these datasets are detailed in Appendix A.1.
In line with the experimental setup described in (Liu et al., 2023b), we only utilized traffic data from 2019. Sensor readings, originally recorded at 5-minute intervals, were aggregated into 15-minute windows, yielding 96 windows per day. Each sub-dataset was chronologically divided into training, validation, and test sets at a ratio of 6:2:2. The benchmark task was to predict the next 12 steps for each sensor at each timestamp. Performance was measured using the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE).
5.2. Baseline Method
Our benchmark experiment incorporates 14 baseline methods. Besides the 11 baseline methods [^3] HL, LSTM (Hochreiter and Schmidhuber, 1997), DCRNN (Li et al., 2017), STGCN (Yu et al., 2017), ASTGCN (Guo et al., 2019), GWNET (Wu et al., 2019), AGCRN (Bai et al., 2020), STGODE (Fang et al., 2021), DSTAGNN (Lan et al., 2022), D2STGNN (Shao et al., 2022d), and DGCRN (Li et al., 2023a) evaluated by Liu et al. (Liu et al., 2023b), we also explore general time series forecasting baseline methods such as the one-nearest-neighbor regressor (1NN), linear model, and TSMixer (Chen et al., 2023). It is important to note that certain baseline methods are not included for the CA dataset due to scalability issues preventing the completion of these methods’ experiments. The details about these methods are included in Appendix A.2.
Table 1. Performance comparisons. We bold the best-performing results. The performance reported in the “Average” column is computed by averaging over 12 predicted time steps. The absence of baselines on the GLA and CA datasets indicates that the models incur an out-of-memory issue. Param: the number of learnable parameters. K: $10^{3}$. M: $10^{6}$.